Good Evening, I'm struggling with this word problem. Can someone Please help?
A moneybox contains only 20-cent coins and 50-cent coins. If there are 57 coins in the moneybox and their total value is $24.30, how many 20-cent coins and how many 50-cent coins are there?

Answer by rothauserc(4718)

we have two equations
x +y = 57
.20x + .50y = 24.30
solve first equation for x, then substitute for x in second equation
x = 57 -y
.20*(57-y) + .50y = 24.30
11.40 -.20y + .50y = 24.30
.30y + 11.40 = 24.30
subtract 11.40 from both sides of =
.30y = 12.90
y = 43
substitute in first equation
x +43 = 57
x = 14
substitute for x and y in second equation
.20*14 +.50*43 = 24.30
2.80 + 21.50 = 24.30
24.30 = 24.30
our answers check
there are 14 ( 20 cent coins ) and 43 ( 50 cent coins )